OpenStudy (anonymous):

Please help? i'll do anything in return!! EMERGENCY!!!!! Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Provide the solution to the system. 5x - y = -1 7x - 3y = 13

hero (hero):

Hint: Isolate y in the first equation, then substitute the expression for y into the second equation: 5x - y = -1 5x + 1 = y

OpenStudy (anonymous):

thannk you :) I got 5 x - -1

hero (hero):

You got what?

OpenStudy (anonymous):

5x -1

OpenStudy (anonymous):

is that right? @Hero

hero (hero):

From where did you get that? Why don't you post your steps please.

OpenStudy (anonymous):

5x - y = -1 5x + 1 = y 5x - 1 because you want to get rid of the addition sign

OpenStudy (anonymous):

5x-y=-1 5x-y+(y)=-1+(y) 5x=-1+y 5x+(1)=-1+y+(1) therefore y=5x+1.............(1)

hero (hero):

@InsanelyChaotic, your reasoning behind using 5x - 1 is flawed.

OpenStudy (anonymous):

try to do the same for the second equation then equate the two y values. u'll find the x value, i think u can help urself now?

OpenStudy (anonymous):

i cant get it.....

OpenStudy (anonymous):

try to isolate y in the second equation the way i did then we'll move bit by bit